OpenStudy (anonymous):

What is the nuclear binding energy in joules for an atom of helium–4? Assume the following: Mass defect = 5.0531 × 10-29 kilograms. Use E = mc2, with c = 3 × 108 m/s.

OpenStudy (aaronq):

\(\sf \huge \underbrace{E}_{\normalsize energy}=\underbrace{m}_{\normalsize mass}*\underbrace{c^2}_{\normalsize speed~of~light}\) \(\sf \large E=(5.0531 × 10^{-29} kg)*(3 × 10^8~ m/s)^2\) Units: \(\large \sf \dfrac{kg*m^2}{s^2}= J ~(joule)\)